### v1

# PSMGD: Periodic Stochastic Multi-Gradient Descent for Fast Multi-Objective Optimization

Abstract
Multi-objective optimization (MOO) lies at the core of many machine learning (ML) applications that involve multiple, potentially conflicting objectives (e.g., multi-task learning, multi-objective reinforcement learning, among many others). Despite the long history of MOO, recent years have witnessed a surge in interest within the ML community in the development of gradient manipulation algorithms for MOO, thanks to the availability of gradient information in many ML problems. However, existing gradient manipulation methods for MOO often suffer from long training times, primarily due to the need for computing dynamic weights by solving an additional optimization problem to determine a common descent direction that can decrease all objectives simultaneously. To address this challenge, we propose a new and efficient algorithm called Periodic Stochastic Multi-Gradient Descent (PSMGD) to accelerate MOO. PSMGD is motivated by the key observation that dynamic weights across objectives exhibit small changes under minor updates over short intervals during the optimization process. Consequently, our PSMGD algorithm is designed to periodically compute these dynamic weights and utilizes them repeatedly, thereby effectively reducing the computational overload. Theoretically, we prove that PSMGD can achieve state-of-the-art convergence rates for strongly-convex, general convex, and non-convex functions. Additionally, we introduce a new computational complexity measure, termed backpropagation complexity, and demonstrate that PSMGD could achieve an objective-independent backpropagation complexity. Through extensive experiments, we verify that PSMGD can provide comparable or superior performance to state-of-the-art MOO algorithms while significantly reducing training time.
Codehttps://github.com/MarcuXu/PSMG
Extended versionhttps://aaai.org/example
1 Introduction
Just as the human world is full of diverse and potentially conflicting goals, many machine learning paradigms are also multi-objective, such as multi-task learning (Caruana 1997; Sener and Koltun 2018), multi-objective reinforcement learning (Hayes et al. 2022), learning-to-rank (Mahapatra et al. 2023; Mahapatra, Dong, and Momma 2023), among others. At the core of these machine learning paradigms lies multi-objective optimization (MOO), which aims to optimize multiple potentially conflicting objectives simultaneously. Formally, MOO can be mathematically cast as:
$\min_{{\mathbf{x}}\in\mathcal{D}}\mathbf{F}({\mathbf{x}}):=[f_{1}({\mathbf{x}}),\cdots,f_{S}({\mathbf{x}})],$ (1)
where ${\mathbf{x}}\in\mathcal{D}\subseteq\mathbb{R}^{d}$ is the model parameter, and $f_{s}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, $s\in[S]$ is the objective function.
The most common approach for handling MOO is to optimize a weighted average of the multiple objectives, also known as static/unitary weighting or linear scalarization (Kurin et al. 2022; Xin et al. 2022). This approach transforms MOO into a single-objective problem, for which there are various off-the-shelf methods available. The linear scalarization approach has a fast per-iteration runtime but suffers from slow convergence and poor performance due to potential conflicts among objectives. To address this problem, there has been a surge of interest in recent years in developing gradient manipulation algorithms for MOO (Kendall, Gal, and Cipolla 2018; Chen et al. 2018; Yu et al. 2020; Chen et al. 2020; Javaloy and Valera 2021; Liu et al. 2021a; Liu and Vicente 2021; Chen et al. 2024; Xiao, Ban, and Ji 2024). The key idea is to calculate dynamic weights to avoid conflicts and find a direction $\mathbf{d}({\mathbf{x}})$ that decreases all objective functions simultaneously. This common descent vector $\mathbf{d}({\mathbf{x}})$ is often determined by solving an additional optimization problem that involves all objective gradients. Updating the model with the common descent direction $\mathbf{d}$ ensures a decrease in every objective. While these approaches show improved performance, the extra optimization process requires a substantial amount of time. This is particularly evident when the number of objectives is large and the model parameters are high-dimensional, which could significantly prolong the MOO training process. Existing empirical evaluations consistently demonstrate that gradient manipulation algorithms typically necessitate much longer training times (Kurin et al. 2022; Liu et al. 2024) (also shown in Sec 4). To this end, a natural question arises: Can we design efficient gradient manipulation algorithms for fast MOO with theoretical guarantees?
In this paper, we present Periodic Stochastic Multi-Gradient Descent (PSMGD), a simple yet effective algorithm to accelerate the MOO process. Our PSMGD is motivated by the key observation that dynamic weights across objectives exhibit small changes under minor updates over short intervals. Thus, our PSMGD algorithm is designed to periodically compute these dynamic weights and utilizes them repeatedly, thereby effectively reducing the computational overload and achieving a faster MOO training process. Our contributions are summarized as follows:
Table 1: Comparison of different algorithms for MOO problem in stochastic first order oracle.
Convexity Algorithm Assume Lipschitz Convergence BP Complexity
continuity of $\bm{\lambda}^{*}({\mathbf{x}})$ Rate
Strongly Convex SMG (Liu and Vicente 2021) $\mathcal{O}(\frac{1}{T})$ $\mathcal{O}(\frac{S}{\epsilon})$
MoDo (Chen et al. 2024) $\mathcal{O}(\frac{1}{T})$ $\mathcal{O}(\frac{S}{\epsilon})$
CR-MOGM (Zhou et al. 2022) $\mathcal{O}(\frac{1}{T})$ $\mathcal{O}(\frac{S}{\epsilon})$
PSMGD $\mathcal{O}(\frac{1}{T})$ $\mathcal{O}(\frac{S}{\epsilon R}+\frac{(R-1)}{\epsilon R})$
General Convex SMG (Liu and Vicente 2021) $\mathcal{O}(\frac{1}{\sqrt{T}})$ $\frac{S}{\epsilon^{2}}$
CR-MOGM (Zhou et al. 2022) $\mathcal{O}(\frac{1}{\sqrt{T}})$ $\frac{S}{\epsilon^{2}}$
PSMGD $\mathcal{O}(\frac{1}{\sqrt{T}})$ $\mathcal{O}(\frac{S}{\epsilon^{2}R}+\frac{(R-1)}{\epsilon^{2}R})$
Non-Convex CR-MOGM (Zhou et al. 2022) $\mathcal{O}(\frac{1}{\sqrt{T}})$ $\frac{S}{\epsilon^{2}}$
MoCO (Fernando et al. 2022) $\mathcal{O}(\frac{1}{\sqrt{T}})$ $\frac{S}{\epsilon^{2}}$
MoDo (Chen et al. 2024) $\mathcal{O}(\frac{1}{\sqrt{T}})$ $\frac{S}{\epsilon^{2}}$
SDMGrad (Xiao, Ban, and Ji 2024) $\mathcal{O}(\frac{1}{\sqrt{T}})$ $\frac{S}{\epsilon^{2}}$
PSMGD $\mathcal{O}(\frac{1}{\sqrt{T}})$ $\mathcal{O}(\frac{S}{\epsilon^{2}R}+\frac{(R-1)}{\epsilon^{2}R})$
  • 1. BP Complexity measures the total number of backpropagation to achieve certain metric ($\epsilon$), which is defined in Def 3.
  • 2. Notations: T is the number of iterations, S is the number of objectives, R is the weight calculation hyper-parameter.
  • Built upon the key observation, we introduce PSMGD, a new and efficient gradient manipulation method algorithm. By calculating dynamic weights across objectives infrequently, PSMGD significantly alleviates computational burden and reduces training time.
  • We conduct a theoretical analysis of PSMGD for strongly convex, general convex, and non-convex functions, demonstrating that PSMGD achieves state-of-the-art convergence rates comparable to existing MOO methods (Table 1). We also introduce a new computational complexity measure, backpropagation (BP) complexity, to quantify computational workload, and further show that PSMGD can achieve an objective-independent backpropagation complexity. Compared with existing MOO methods, this shows a superior BP complexity with the linear speedup with respect to the number of objectives.
  • Through comprehensive evaluation, PSMGD shows comparable or even superior performance compared to existing MOO methods, with significantly less training time.
2 Related works
The MOO problem, as stated in Eq. 1, has a long and rich history that dates back to the 1950s. MOO algorithms can be broadly categorized into two main groups. The first category comprises gradient-free methods, such as evolutionary MOO algorithms and Bayesian MOO algorithms (Zhang and Li 2007; Deb et al. 2002; Belakaria et al. 2020; Laumanns and Ocenasek 2002). These methods are more suitable for small-scale problems but are less practical for high-dimensional MOO models, such as deep neural networks. The second category is the gradient-based approach by utilizing (stochastic) gradients (Fliege and Svaiter 2000; Désidéri 2012; Fliege, Vaz, and Vicente 2019; Liu and Vicente 2021), making them more practical for high-dimensional MOO problems. In this work, we primarily focus on gradient-based approaches to solve MOO in high-dimensional deep-learning models. In what follows, we provide a brief overview of two typical approaches for gradient-based methods for MOO: linear scalarization (LS) and gradient manipulation methods.
1) LS: A straightforward approach for MOO 1 is to transform it into a single objective problem by using pre-defined weights $\{\lambda_{s},s\in[s]\}$: $\min_{{\mathbf{x}}\in\mathcal{D}}\sum_{s\in[S]}\lambda_{s}f_{s}({\mathbf{x}}).$ Thanks to the LS, one can leverage many existing and familiar single-objective methods, such as stochastic gradient descent, for training at each iteration. However, in LS, it is not uncommon to see conflicts among multiple objectives during the optimization process, i.e., $\nabla f_{s}({\mathbf{x}}),\nabla f_{s^{\prime}}({\mathbf{x}})\right><0$. This implies that the update using static weights may decrease some objectives, while inevitably increasing others at the same time and leading to slow convergence and poor performance.
2) Gradient Manipulation Methods: A popular alternative is to dynamically weight gradients across objectives to avoid such conflicts and obtain a direction to decrease all objectives simultaneously. For example, multiple gradient descent algorithm (MGDA) (Fliege and Svaiter 2000) is seeking to find the common descent direction by solving the following optimization problem given each objective’s gradient: $(\mathbf{d},\beta)\in\operatorname*{arg\,min}_{\mathbf{d}\in\mathbb{R}^{S},\beta\in\mathbb{R}}\beta+\frac{1}{2}\|\mathbf{d}\|^{2},s.t.,\nabla f_{s}({\mathbf{x}})^{T}\mathbf{d}-\beta\leq 0,\forall s\in[S].$ By doing so, if ${\mathbf{x}}$ represents a first-order Pareto stationary point, then $(\mathbf{d},\beta)=(\mathbf{0},0)$. Otherwise, we can find the $\mathbf{d}$ such that $\nabla f_{s}({\mathbf{x}})^{T}\mathbf{d}\leq\beta<0,s\in[S]$ as the solution. Using such a non-conflicting direction $\mathbf{d}$ to update the model, i.e., ${\mathbf{x}}\leftarrow{\mathbf{x}}+\eta{\mathrm{d}}$ where $\eta$ is the learning rate, has been shown to achieve better performance in practice (Sener and Koltun 2018; Liu et al. 2021a). Following this token, many gradient manipulation methods attempt to obtain a common descent direction through various formulations and solutions. For example, GradDrop (Chen et al. 2020) randomly dropped out highly conflicted gradients, RotoGrad (Javaloy and Valera 2021) rotated objective gradients to alleviate the conflict, and many other methods exploring similar principles (Kendall, Gal, and Cipolla 2018; Chen et al. 2018; Yu et al. 2020; Liu et al. 2021a; Liu and Vicente 2021; Chen et al. 2024; Xiao, Ban, and Ji 2024). However, due to the additional optimization process required to generate such $\mathbf{d}$, gradient manipulation methods often incur an expensive per-iteration cost, thereby prolonging training time. In this work, we propose a new algorithm, PSMGD, which combines the advantages of both approaches and offers a fast per-iteration runtime, a fast overall convergence speed, and improved model performance.
3 Periodic stochastic multi-gradient descent
In this section, we first present the basic concept of MOO, followed by a pedagogical example comparing linear scalarization with MGDA, a representative of gradient manipulation methods. Based on one observation of the stable variation of dynamic weights for MGDA, we propose the PSMGD algorithm, followed by its convergence analyses.
3.1 Preliminaries of MOO
Analogous to the stationary and optimal solutions in single-objective optimization, MOO seeks to adopt the notion of Pareto optimality/stationarity:
Definition 1 ((Weak) Pareto Optimality).
For any two solutions ${\mathbf{x}}$ and ${\mathbf{y}}$, we say ${\mathbf{x}}$ dominates ${\mathbf{y}}$ if and only if $f_{s}({\mathbf{x}})\leq f_{s}({\mathbf{y}}),\forall s\in[S]$ and $f_{s}({\mathbf{x}})<f_{s}({\mathbf{y}}),\exists s\in[S]$. A solution ${\mathbf{x}}$ is Pareto optimal if it is not dominated by any other solution. One solution ${\mathbf{x}}$ is weakly Pareto optimal if there does not exist a solution ${\mathbf{y}}$ such that $f_{s}({\mathbf{x}})>f_{s}({\mathbf{y}}),\forall s\in[S]$.
Definition 2 (Pareto Stationarity).
A solution ${\mathbf{x}}$ is said to be Pareto stationary if there is no common descent direction $\mathbf{d}\in\mathbb{R}^{d}$ such that $\nabla f_{s}({\mathbf{x}})^{\top}\mathbf{d}<0,\forall s\in[S]$.
Similar to solving single-objective non-convex optimization problems, finding a Pareto-optimal solution in MOO is NP-Hard in general. As a result, it is often of practical interest to find a solution satisfying Pareto-stationarity (a necessary condition for Pareto optimality). Following Definition 2, if ${\mathbf{x}}$ is not a Pareto stationary point, we can find a common descent direction $\mathbf{d}\in\mathbb{R}^{d}$ to decrease all objectives simultaneously, i.e., $\nabla f_{s}({\mathbf{x}})^{\top}\mathbf{d}<0,\forall s\in[S]$. If no such a common descent direction exists at ${\mathbf{x}}$, then ${\mathbf{x}}$ is a Pareto stationary solution. For example, multiple gradient descent algorithm (MGDA) (Désidéri 2012) searches for an optimal weight $\bm{\lambda}_{*}$ of gradients $\nabla\mathbf{F}({\mathbf{x}}):=\{\nabla f_{s}({\mathbf{x}}),\forall s\in[S]\}$ by solving $\bm{\lambda}_{*}({\mathbf{x}})=\operatorname*{argmin}_{\bm{\lambda}}\|\bm{\lambda}^{\top}\nabla\mathbf{F}({\mathbf{x}})\|^{2}$, which is the dual of the original problem. Then, a common descent direction can be chosen as: $\mathbf{d}=\bm{\lambda}_{*}^{\top}\nabla\mathbf{F}({\mathbf{x}})$. MGDA performs the iterative update rule: ${\mathbf{x}}\leftarrow{\mathbf{x}}-\eta\mathbf{d}$ until a Pareto optimal/stationary point is reached, where $\eta$ is a learning rate. Many gradient manipulation algorithms have been inspired by MGDA. In the next subsection, we compare two typical gradient-based approaches using a pedagogical example, with MGDA representing the gradient manipulation approach.
Figure 1: The convergence behaviors on a pedagogical example with 10 runs for each algorithm: (a) Solutions obtained from random linear scalarization. (b) Solutions obtained from the MGDA method. (c) Solutions obtained from the PSMGD method proposed in this paper, with weights updated every 4 iterations ($R=4$) through the training. (d) We visualize weights changing curves over iterations in 5 runs. This method successfully generates a set of widely distributed Pareto solutions with different trade-offs. Details of the pedagogical example can be found in Section 3.2.
3.2 A pedagogical example
We consider the Fonseca problem(Fonseca and Fleming 1996; Pardalos et al. 2017), which is to solve a two-objective optimization problem, $\min_{{\mathbf{x}}}[f_{1}({\mathbf{x}}),f_{2}({\mathbf{x}})]$. These two objective functions are defined by $f_{1}({\mathbf{x}})=1-\exp\left(-\sum_{i=1}^{d}\left(x_{d}-\frac{1}{\sqrt{d}}\right)^{2}\right),f_{2}({\mathbf{x}})=1-\exp\left(-\sum_{i=1}^{d}\left(x_{d}+\frac{1}{\sqrt{d}}\right)^{2}\right),$ where ${\mathbf{x}}=(x_{1},x_{2},...,x_{d})$ representing the d-dimensional decision variable. The problem exhibits a non-convex Pareto front in the objective space. We run three algorithms on this pedagogical example: linear scalarization, MGDA, and our proposed PSMGD.
Setting. We run each algorithm 10 times. In Figure 1, we show the trajectory of different methods from random starting points. The blue dashed lines represent the trajectories, with the final solutions highlighted by red dots.
Observation 1: Linear scalarization can only converge to specific points and is unable to explore different trade-offs among objectives. we run the linear scalarization method by using randomly static weights. As shown in Figure 1(a), it is obvious that linear scalarization can only converge to two specific points (see red dots), with either $f_{1}({\mathbf{x}})$ or $f_{2}({\mathbf{x}})$ being under-optimized. In other words, linear scalarization cannot fully explore the Pareto front, which is consistent with previous results (Miettinen 1999; Pardalos et al. 2017). In contrast, MGDA can find trade-off points, and our proposed PSMGD successfully identifies a set of well-distributed Pareto solutions with different trade-offs, shown in Figure 1(b,c). We note that linear scalarization usually runs faster for MOO in each iteration than gradient manipulation methods due to its simplicity, as shown in many existing works (Kurin et al. 2022; Liu et al. 2024) and confirmed by our experimental results (see Sec 4). However, our observation highlights the necessity of gradient manipulation methods and motivates us to develop faster gradient manipulation techniques for MOO.
Observation 2: The dynamic weights calculated by MGDA converge quickly and exhibit stability. As shown in Figure 1(d), we visualize the dynamic weights for the Fonseca problem. In this simple problem, the weights converge within 5 iterations, after which they remain unchanged. In more complex problems, such as multi-task learning using deep learning (see Sec B in Appendix), the dynamic weights calculated by gradient manipulation methods also demonstrate stability. We verified stability by examining two scales, epoch and iteration. The weights show consistent behavior with a stable average value and low variability across both scales, epoch and iteration, as depicted in Figure 6(a) and 6(b). This observation suggests a new approach to develop more efficient MOO algorithms. Specifically, we can calculate and update the dynamic weights periodically in a more lazy and thus efficient manner.
3.3 PSMGD
Based on our observations, we propose a new and efficient algorithm called Periodic Stochastic Multi-Gradient Descent (PSMGD) for MOO, as shown in Algorithm 1. Our PSMGD algorithm is designed to periodically compute these dynamic weights and utilizes them repeatedly, thereby reducing the computational load effectively. Specifically, in each iteration $t\in[T]$, we offer two options: 1) If $t\%R==0$, we calculate the dynamic weights $\hat{\bm{\lambda}}_{t}^{*}$ based on the stochastic gradient of each objective $\nabla f_{s}({\mathbf{x}}_{t},\xi_{t}),s\in[S]$, by solving optimization problem  2. Following this, we apply momentum to $\bm{\lambda}$ to further stabilize the weights: $\bm{\lambda}_{t}=\alpha_{t}\bm{\lambda}_{t-R}+(1-\alpha_{t})\hat{\bm{\lambda}}_{t}^{*}$. 2) Otherwise, we reuse the dynamic weights: $\bm{\lambda}_{t}=\bm{\lambda}_{t-1}.$ With predefined $\bm{\lambda}_{t}$, it is worth pointing out that the MOO can be naturally transformed into a single-objective problem by using pre-defined weighted sum of the objectives, thereby requiring only one backpropagation step. Subsequently, after obtaining the weights across objectives, we can approximate the common descent direction by $\mathbf{d}_{t}=\bm{\lambda}_{t}^{T}\nabla\mathbf{F}({\mathbf{x}}_{t},\xi_{t})$. Then the model can be updated by ${\mathbf{x}}_{t+1}={\mathbf{x}}_{t}-\eta_{t}\mathbf{d}_{t}$.
Algorithm 1 Periodic Stochastic Multi-Gradient Descent (PSMGD)
1:  Initialize model parameter ${\mathbf{x}}_{0}$, learning rate $\eta$, and hyper-parameter $R$.
2:  for $t=0,\dots,T-1$ do
[…]
4:         Compute $\hat{\bm{\lambda}}_{t}^{*}\in[0,1]^{S}$ by solving
$\min_{\bm{\lambda}}\Big{\|}\sum\nolimits_{s\in[S]}\lambda_{s}\nabla f_{s}({\mathbf{x}}_{t},\xi_{t})\Big{\|}^{2},$
[…]
5:         Update: $\bm{\lambda}_{t}=\alpha_{t}\bm{\lambda}_{t-R}+(1-\alpha_{t})\hat{\bm{\lambda}}_{t}^{*}.$
6:     Otherwise: $\bm{\lambda}_{t}=\bm{\lambda}_{t-1}.$blacktriangleright Reuse the weights
7:     Update the model: ${\mathbf{x}}_{t+1}={\mathbf{x}}_{t}-\eta_{t}\mathbf{d}_{t}$, where $\mathbf{d}_{t}=\sum\nolimits_{s\in[S]}\lambda_{t,s}\nabla f_{s}({\mathbf{x}}_{t},\xi_{t})$.
8:  end for
3.4 Convergence analysis
Before providing the convergence analysis, we have the following assumptions:
Assumption 3.1 (L-Lipschitz continuous).
There exists a constant $L>0$ such that $\|\nabla f_{s}({\mathbf{x}})-\nabla f_{s}({\mathbf{y}})\|\leq L\|{\mathbf{x}}-{\mathbf{y}}\|,\forall{\mathbf{x}},{\mathbf{y}}\in\mathbb{R}^{d},s\in[S]$.
Assumption 3.2 (Unbiased stochastic gradient with bounded variance).
We assume the stochastic gradient estimation of each objective function is unbiased with bounded variance.
$\mathbb{E}\left[\nabla f_{s}({\mathbf{x}},\xi)\right]=\nabla f_{s}({\mathbf{x}}),$ (3)
$\mathbb{E}\left[\|\nabla f_{s}({\mathbf{x}},\xi)-\nabla f_{s}({\mathbf{x}})\|^{2}\right]\leq\sigma^{2},s\in[S].$ (4)
Assumption 3.3 (Bounded Weights).
There exists a constant $B$ s.t., $0\leq\lambda_{t,s}\leq B,\sum_{s\in[S]}\lambda_{t,s}\geq 1,\forall s\in[S],t\in[T]$.
Assumption 3.4 (Bounded Gradient).
The gradient of each objective is bounded, i.e., there exists a constant $H>0$ such that $\|\nabla f_{s}({\mathbf{x}})\|\leq H,\forall s\in[S]$.
These four assumptions are widely-used assumptions in MOO (Liu and Vicente 2021; Zhou et al. 2022). For notation clarity, we have the following definition: $G({\mathbf{x}}_{t},\bm{\lambda}_{t})=\sum_{s\in[S]}\lambda_{t,s}f_{s}({\mathbf{x}}_{t}),\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})=\sum_{s\in[S]}\lambda_{t,s}\nabla f_{s}({\mathbf{x}}_{t})$. With these assumptions and definitions, we have the following convergence rates:
Theorem 3.5 (Non-Convex Functions).
Under Assumptions 3.13.4, when each objective is bounded by $F$ ($f_{s}({\mathbf{x}})\leq F,s\in[S]$), the sequence of iterates generated by the PSMGD Algorithm in non-convex functions satisfies:
$\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]$ (5)
$\leq\frac{4SBF}{T\eta_{T}}+\frac{2F}{T}\sum_{\begin{subarray}{c}t\%R==0\\ t\neq 0\end{subarray}}\frac{\mathbb{E}\left[\sum_{s\in[S]}(1-\alpha_{t})|\hat{\lambda}_{t}^{s}-\lambda_{t-R}^{s}|\right]}{\eta_{t}}$
$+2LS^{2}B^{2}\sigma^{2}\frac{1}{T}\sum_{t\in[T]}\eta_{t}+\frac{4S^{3/2}B^{2}H\sigma}{T}\sum_{t\%R==0}(1-\alpha_{t}).$
Setting $1-\alpha_{t}=\min\{\frac{\eta_{t}}{\eta_{1}},\frac{\eta_{t}}{\eta_{1}\sqrt{t}\max_{s\in[S]}|\hat{\lambda}_{t}^{s}-\lambda_{t-R}^{s}|}\},\eta_{t}=\mathcal{O}(\frac{1}{\sqrt{t}})$, the convergence rate is $\mathcal{O}(\frac{1}{\sqrt{T}}).$
Theorem 3.6 (General Convex Functions).
Under Assumptions 3.13.4, when the distance from sequence to Pareto set is bounded ($\|{\mathbf{x}}_{t}-{\mathbf{x}}_{*}\|\leq D$), the sequence of iterates generated by the PSMGD in general convex functions satisfies:
$\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[G({\mathbf{x}}_{t},\bm{\lambda}_{t})-G({\mathbf{x}}_{*},\bm{\lambda}_{t})]$ (6)
$\leq\frac{\|{\mathbf{x}}_{1}-{\mathbf{x}}_{*}\|^{2}}{T\eta_{1}}+2D\sigma S^{3/2}B\frac{1}{T}\sum_{t\%R==0}(1-\alpha_{t})$
$\quad+(2S^{2}B^{2}\sigma^{2}+2S^{2}B^{2}H^{2})\frac{1}{T}\sum_{t\in[T]}\eta_{t}.$
Setting $\eta_{t}=\frac{1}{\sqrt{t}}$ and $(1-\alpha_{t})=\eta_{t}$, the convergence rate is $\mathcal{O}(1/\sqrt{T}).$
Theorem 3.7 ($\mu$-Strongly Convex Functions).
Under Assumptions 3.13.4 and we further assume each objective function $f_{s}({\mathbf{x}}),s\in[S]$ is $\mu-$strongly convex and each objective is bounded by $F$ ($f_{s}({\mathbf{x}})\leq F,s\in[S]$), the sequence of iterates generated by the PSMGD satisfies:
$\mathbb{E}[G({\mathbf{x}}_{t+1},\bm{\lambda}_{t+1})-G({\mathbf{x}}_{t+1}^{*},\bm{\lambda}_{t+1})]$
$\leq(1-2\mu\eta_{t})\mathbb{E}[G({\mathbf{x}}_{t},\bm{\lambda}_{t})-G({\mathbf{x}}_{t}^{*},\bm{\lambda}_{t})]+\eta_{t}^{2}\Phi,$ (7)
where $\Phi=2LS^{2}B^{2}\sigma^{2}+4FSB\mathbf{1}\{(t+1)\%R=0\}+4S^{3/2}B^{2}H\sigma\mathbf{1}\{t\%R=0\}$ and $\mathbf{1}$ is the indicator function. Set $\eta=\frac{c}{T}$ with $c>\frac{1}{\mu}$, the convergence rate is
$\mathbb{E}[G({\mathbf{x}}_{T},\bm{\lambda}_{T})-G({\mathbf{x}}_{T}^{*},\bm{\lambda}_{T})]$
$\leq\frac{\max\{2c^{2}\Phi^{{}^{\prime}}(2\mu c-1)^{-1},G({\mathbf{x}}_{0},\bm{\lambda}_{0})-G({\mathbf{x}}_{0}^{*},\bm{\lambda}_{0})\}}{T}$
$=\mathcal{O}(1/T),$ (8)
where $\Phi^{{}^{\prime}}=2LS^{2}B^{2}\sigma^{2}+4FSB/R+4S^{3/2}B^{2}H\sigma/R.$
Remark 3.8.
1) Convergence Metrics. Different convergence metrics are used in existing studies, and we follow convergence conditions outlined in (Tanabe, Fukuda, and Yamashita 2019; Zhou et al. 2022). In the appendix, we provide an explanation of convergence metrics and a comparison with other works. 2) Convergence Rates. With proper hyperparameters, our algorithm, PSMGD , can achieve $\mathcal{O}(1/T)$ for strongly convex functions and $\mathcal{O}(1/\sqrt{T})$ for general convex and non-convex functions. These convergence rates match the state-of-the-art rates in existing MOO algorithms.
The convergence rate reflects the training speed in terms of iterations but does not fully capture the total computational complexity. Similar to how sample complexity is used in single-objective learning, we propose a new metric, Backpropagation (BP) complexity, to quantify the computational workload tailored to MOO in first-order oracle.
Definition 3 (Backpropogation complexity).
We define Backpropagation Complexity as the total number of backpropagation operations required by an algorithm to achieve a specified performance threshold, denoted as $\epsilon$.
Remark 3.9.
Our PSMGD algorithm can achieve BP complexity of $\mathcal{O}(\frac{S}{\epsilon R}+\frac{(R-1)}{\epsilon R})$ for stongly-convex functions and $\mathcal{O}(\frac{S}{\epsilon^{2}R}+\frac{(R-1)}{\epsilon^{2}R})$ for general convex and non-convex functions. Compared to existing MOO methods, PSMGD can achieve a linear speedup in terms of $R$.
Remark 3.10.
If $R=\Omega(S)$, PSMGD exhibits an objective- independent BP complexity, i.e., $\mathcal{O}(\frac{1}{\epsilon})$ for stongly-convex functions and $\mathcal{O}(\frac{1}{\epsilon^{2}})$ for general convex and non-convex functions. These rates indicate that PSMGD requires the same order of computations for MOO as classic SGD does for single-objective learning. We provide a comparison of these rates and complexities in Table 1, which are also validated by extensive experiments in Sec 4.
4 Experiment
In this section, we conduct a comprehensive empirical evaluation of the proposed PSMGD algorithm, focusing primarily on multi-task learning with deep neural networks. The primary goal is to verify the following key points:
  • Improved Performance: Can PSMGD provide improved model performance?
  • Fast Training: Can PSMGD have a fast training process?
  • Ablation Study: How does the hyper-parameter impact the training? (Results In Appendix)
Baseline Algorithms. To conduct an extensive comparison, we use 14 algorithms as the baselines, including a single objective learning baseline, linear scalarization, and 12 advanced MOO algorithms. They are: (1) Single task learning (STL), training independent models $f_{S}({\mathbf{x}})$ for all objectives; (2) Linear scalarization (LS) baseline; (3) Scale-invariant (SI) that minimizes $\sum_{s\in[S]}\log f_{s}({\mathbf{x}})$; (4) Dynamic Weight Average (DWA(Liu, Johns, and Davison 2019) adaptively updates weights based on the comparative rate of loss reduction for each objective; (5) Uncertainty Weighting (UW(Kendall, Gal, and Cipolla 2018) leverages estimated task uncertainty to guide weight assignments; (6) Random Loss Weighting (RLW(Lin, Feiyang, and Zhang 2021) samples objective weighting with log-probabilities by normal distribution; (7) MGDA (Sener and Koltun 2018) identifies a joint descent direction that concurrently decreases all objectives; (8) PCGrad (Yu et al. 2020) projects each objective gradient onto the normal plan of the gradients to avoid conflicts; (9) CAGrad (Liu et al. 2021a) balances the average loss while guaranteeing a controlled minimum improvement for each objective; (10) IMTL-G (Liu et al. 2021b) determines the update direction by having equal projections on objective gradients; (11) GradDrop (Chen et al. 2020) randomly drops out certain dimensions of the objective gradients for their level of conflict; (12) NashMTL (Navon et al. 2022) determines the game’s solution that is advantageous for all goals by a bargaining game; (13) FAMO (Liu et al. 2024) is a dynamic weighting method that reduces objective losses in space and time in a balanced manner; (14) FairGrad (Maheshwari and Perrot 2022) aims to achieve group fairness in MOO by dynamically re-weighting; (15) SDMGrad (Xiao, Ban, and Ji 2024) utilizes direction-oriented regularization.
Datasets. We evaluate on 5 multi-task learning datasets with objective/task numbers ranging from 2 to 40, covering scenarios in regression, classification, and dense prediction. For regression, we choose QM-9 dataset (Blum and Reymond 2009) (11 tasks), a widely used benchmark in graph neural network learning. For image classification, we use Multi-MNIST (Sener and Koltun 2018) (2 tasks) and CelebA dataset (Liu et al. 2015) (40 tasks). For dense prediction, CityScapes  (Cordts et al. 2016) (2 tasks) and NYU-v2 (Silberman et al. 2012) (3 tasks) are used in our experiments. NYU-v2 is a dataset of indoor scenes with 1449 RGBD images and dense per-pixel labeling across 13 classes. CityScapes is similar to NYU-v2 but boasts 5K street-view RGBD images with per-pixel annotations. We present the results for QM-9 and NYU-v2 in this section, with the remaining results and experimental settings detailed in Appendix B.
Method $\mu$ $\alpha$ $\epsilon_{\text{HOMO}}$ $\epsilon_{\text{LUMO}}$ $\langle R^{2}\rangle$ ZPVE $U_{0}$ $U$ $H$ $G$ $c_{v}$ MR $\downarrow$ $\Delta m\%$downarrow
MAE $\downarrow$
STL 0.07 0.18 60.6 53.9 0.50 4.53 58.8 64.2 63.8 66.2 0.07
LS 0.11 0.33 73.6 89.7 5.20 14.06 143.4 144.2 144.6 140.3 0.13 9.00 177.6
SI 0.31 0.35 149.8 135.7 1.00 4.51 55.3 55.8 55.8 55.3 0.11 5.36 77.8
RLW 0.11 0.34 76.9 92.8 5.87 15.47 156.3 157.1 157.6 153.0 0.14 10.36 203.8
DWA 0.11 0.33 74.1 90.6 5.09 13.99 142.3 143.0 143.4 139.3 0.13 8.64 175.3
UW 0.39 0.43 166.2 155.8 1.07 4.99 66.4 66.8 66.8 66.2 0.12 6.64 108.0
MGDA 0.22 0.37 126.8 104.6 3.23 5.69 88.4 89.4 89.3 88.0 0.12 8.36 120.5
PCGrad 0.11 0.29 75.9 88.3 3.94 9.15 116.4 116.8 117.2 114.5 0.11 7.18 125.7
CAGrad 0.12 0.32 83.5 94.8 3.22 6.93 114.0 114.3 114.5 112.3 0.12 8.18 112.8
IMTL-G 0.14 0.29 98.3 93.9 1.75 5.70 101.4 102.4 102.0 100.1 0.10 6.64 77.2
NashMTL 0.10 0.25 82.9 81.9 2.43 5.38 74.5 75.0 75.1 74.2 0.09 3.82 62.0
FAMO 0.15 0.30 94.0 95.2 1.63 4.95 70.82 71.2 71.2 70.3 0.10 5.09 58.5
FairGrad 0.12 0.25 87.57 84.00 2.15 5.07 70.89 71.17 71.21 70.88 0.10 4.09 57.9
PSMGD 0.12 0.25 77.2 74.4 3.01 6.61 103.0 103.5 103.7 101.6 0.09 5.55 92.4
Table 2: Results obtained on the QM-9 dataset. Each experiment is conducted using 3 random seeds, and the mean value is presented. The best average result is highlighted in bold.
Metrics. We have two types of metrics to measure MOO algorithms. I. Model performance metrics. We consider two widely-used metrics to represent the overall performance of one MOO method $m$: (1) $\Delta m\%$, the average per-task performance drop of a method $m$ relative to the STL baseline denoted as $B$: $\bm{\Delta}\bm{m}\%=\frac{1}{N}\sum_{n=1}^{N}(-1)^{\delta_{n}}\frac{(M_{m,n}-M_{B,n})}{M_{B,n}}\times 100,$ where $M_{B,n}$ and $M_{m,n}$ represent the STL and $m$’s value for metric $M_{n}$, respectively. Here, $\delta_{n}$ equals 1 if a higher value of $M_{n}$ is better or 0 if a lower value of $M_{n}$ is better. (2) Mean Rank (MR), the average rank of each method in the tasks. For example, if a method ranks first in every task (whether higher or lower is better), MR will be 1. Note that in practice, lower values of $\Delta m\%$ and MR indicate better overall performance in the final result. II. Convergence metrics. We calculate the training time per epoch and total training time to evaluate the convergence performance of MOO algorithms. Based on various datasets, we also calculate the test loss for a visualization comparison through the training process.
4.1 Regression: QM-9
QM-9 dataset (Blum and Reymond 2009) is a crucial benchmark widely used in the field of graph neural network learning, especially in chemical informatics applications. With over 13K molecules, each molecule is intricately depicted as a graph, with nodes representing atoms and edges symbolizing the chemical bonds between them. These graphs feature detailed node and edge characteristics such as atomic number, atom type, bond type, and bond order. The goal is to predict 11 molecule properties. Utilizing 110K molecules from QM9 in PyTorch Geometric (Fey and Lenssen 2019) for training, 10K for validation, and the remaining 10K for testing.
Performance. We assess two common metrics, MR and $\Delta m\%$, for various MOO algorithms to ensure effective evaluations. Table 2 shows that PSMGD achieves comparable or even better performance in both MR and $\Delta m\%$ when compared with existing baselines. Specifically, it has the best performance in 3 out of the 11 total tasks, namely $\alpha$, $\epsilon_{\text{LUMO}}$, and $c_{v}$. When we take a closer look at the training process, we observe that PSMGD has a faster and more smooth process as illustrated in Figure 2. Specifically, PSMGD demonstrates a smoother descent towards lower test loss at least in the early stages. In addition, PSMGD exhibits resistance to overfitting, distinguishing itself from IMTL-G and NashMTL due to its straightforward yet efficient structure.
Convergence measurements. In Table 3, we show the training time per epoch and the total training time required to achieve a certain loss. In Figure 2, we use the average Mean Absolute Error (MAE) across all molecular property prediction tasks as test loss, calculated after scaling both the predictions and the ground truth values by the standard deviation. PSMGD ranks third in training speed per epoch but achieves the average test loss (<100, <75) in the shortest overall time. The visualization results can be found in Appendix B.1.
Method Training time $\downarrow$ Avg loss $\downarrow$ Avg loss $\downarrow$
(mean per epoch) < 100 < 75
LS 1.70 75.68 170.28
MGDA 12.94 436.92 754.68
PCGrad 6.15 259.56 630.36
CAGrad 5.29 233.64 451.35
IMTL-G 5.38 275.40 604.80
FAMO 1.96 72.52 131.32
NashMTL 8.45 287.30 371.80
FairGrad 5.40 162.60 314.36
PSMGD 2.23 69.13 109.27
Table 3: Results obtained on the QM-9 dataset. Each experiment is conducted using 3 random seeds, and the mean value is presented. The best average result is highlighted in bold.
Segmentation Depth Surface Normal
Method mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ Angle Dist $\downarrow$ Within $t^{\circ}$uparrow MR $\downarrow$ $\Delta m\%$downarrow
Mean Median 11.25 22.5 30
STL 38.30 63.76 0.6754 0.2780 25.01 19.21 30.14 57.20 69.15
LS 39.29 65.33 0.5493 0.2263 28.15 23.96 22.09 47.50 61.08 11.44 5.59
SI 38.45 64.27 0.5354 0.2201 27.60 23.37 22.53 48.57 62.32 10.11 4.39
RLW 37.17 63.77 0.5759 0.2410 28.27 24.18 22.26 47.05 60.62 14.11 7.78
DWA 39.11 65.31 0.5510 0.2285 27.61 23.18 24.17 50.18 62.39 10.44 3.57
UW 36.87 63.17 0.5446 0.2260 27.04 22.61 23.54 49.05 63.65 10.11 4.05
MGDA 30.47 59.90 0.6070 0.2555 24.88 19.45 29.18 56.88 69.36 8.11 1.38
PCGrad 38.06 64.64 0.5550 0.2325 27.41 22.80 23.86 49.83 63.14 10.67 3.97
GradDrop 39.39 65.12 0.5455 0.2279 27.48 22.96 23.38 49.44 62.87 9.56 3.58
CAGrad 39.79 65.49 0.5486 0.2250 26.31 21.58 25.61 52.36 65.58 7.00 0.20
IMTL-G 39.35 65.60 0.5426 0.2256 26.02 21.19 26.20 53.13 66.24 6.33 -0.76
NashMTL 40.13 65.93 0.5261 0.2171 25.26 20.08 28.40 55.47 68.15 4.22 -4.04
FAMO 38.88 64.90 0.5474 0.2194 25.06 19.57 29.21 56.61 68.98 5.11 -4.10
FairGrad 39.74 66.01 0.5377 0.2236 24.84 19.60 29.26 56.58 69.16 3.22 -4.66
SDMGrad 40.47 65.90 0.5225 0.2084 25.07 19.99 28.54 55.74 68.53 3.44 -4.84
PSMGD 35.44 63.78 0.5494 0.2369 24.83 18.89 30.68 58.00 69.84 6.11 -3.62
Table 4: Results obtained on NYU-v2. Each experiment is conducted using 3 random seeds, and the mean value is presented. The best average result is highlighted in bold.
Method Training time $\downarrow$ Semantic loss $\downarrow$ Semantic loss $\downarrow$ Depth loss $\downarrow$ Depth loss $\downarrow$ Normal loss $\downarrow$ Normal loss $\downarrow$
(mean per epoch) < 1.40 < 1.20 < 0.65 < 0.55 < 0.20 < 0.16
LS 1.36 32.14 66.28 46.18 159.36 30.84 152.76
MGDA 3.34 166.43 384.32 218.88 590.96 56.44 249.77
PCGrad 3.31 72.38 154.63 98.79 299.39 75.67 332.29
CAGrad 3.53 63.18 137.89 94.77 351.06 66.69 213.11
IMTL-G 4.63 101.42 239.72 147.52 470.22 87.59 318.09
FAMO 1.65 32.69 68.46 73.35 156.26 34.23 114.13
NashMTL 3.49 76.34 159.62 79.81 277.60 65.93 201.26
FairGrad 2.93 141.82 328.86 148.47 334.53 48.20 177.41
SDMGrad 1.98 31.52 72.89 63.04 200.94 35.46 128.05
PSMGD 1.85 29.28 64.54 43.07 176.66 31.11 113.46
Table 5: Training time per epoch [Min.] and convergence process (averaged over 3 random seeds) in all tasks on NYU-v2.
4.2 Dense Prediction: NYU-v2
NYU-v2 (Silberman et al. 2012) is an indoor scene dataset containing 1449 RGBD images, each annotated with dense per-pixel labeling across 13 distinct classes. The dataset supports 3 tasks for MOO experiments: image segmentation, depth prediction, and surface normal prediction.
Performance. As shown in Table 4, PSMGD achieves better performance in both MR and $\Delta m\%$ compared to existing MOO algorithms. It performs the best in 5 out of the 9 total tasks. In Figure 3, PSMGD achieves consistently low test losses across all tasks compared to other methods in the initial stage, indicating superior performance across all 3 tasks throughout the 200 training epochs.
Convergence measurements. In Figure 3, PSMGD outperforms other MOO methods in reducing test loss for all three tasks on the NYU-v2 dataset, particularly in the early training stages. Table 5 further highlights PSMGD’s efficiency with its third lowest training time per epoch (1.85), significantly outperforming other classical gradient manipulation methods like IMTL-G (Liu et al. 2021b) and NashMTL (Navon et al. 2022). PSMGD achieves the fastest training times by reaching semantic loss in 29.28 and 64.54 minutes, depth loss in 43.07 minutes, and normal loss in 113.46 minutes, when considering the following thresholds: <1.40 and <1.20 for semantic loss, <0.65 for depth loss, and <0.16 for normal loss. The visualization results can be found in Appendix B.1.
5 Conclusion
In this paper, we propose a novel and efficient algorithm, Periodic Stochastic Multi-Gradient Descent (PSMGD), to accelerate MOO. Our PSMGD algorithm periodically computes dynamic weights and reuses them, significantly reducing the computational load and speeding up MOO training. We establish that PSMGD achieves state-of-the-art convergence rates for strongly convex, general convex, and non-convex functions. Moreover, we demonstrate the superior backpropagation (BP) complexity of our PSMGD algorithm. Extensive experiments confirm that PSMGD delivers performance comparable to or better than existing MOO algorithms, with a substantial reduction in training time. We believe future work could explore preference-based solutions or the entire Pareto set using our efficient PSMGD algorithm.
Appendix
Appendix A Proof
A.1 Convergence Analysis
Notation. For notation clarity, we have the following definition: $\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})=\sum_{s\in[S]}\lambda_{t,s}\nabla f_{s}({\mathbf{x}}_{t})$, ^$\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})=\sum_{s\in[S]}\lambda_{t,s}\nabla f_{s}({\mathbf{x}}_{t},\xi_{t})$. $\hat{\bm{\lambda}}_{t}^{*}$ is calculated by QP using stochastic gradient $\nabla f_{s}({\mathbf{x}}_{t},\xi_{t})$, and $\bm{\lambda}_{t}=\alpha_{t}\bm{\lambda}_{t-R}+(1-\alpha_{t})\hat{\bm{\lambda}}_{t}^{*}.$
Convergence Metrics. We note that existing works use similar but slightly different metrics to measure the convergence of MOO. For strongly-convex functions, (Liu and Vicente 2021) uses $\min_{t=1,\dots,T}\sum_{s\in[S]}\left[\hat{\lambda}_{t,s}^{*}f_{s}({\mathbf{x}}_{t})-\bar{\lambda}_{T}f_{s}({\mathbf{x}}_{*})\right]$ where $\bar{\lambda}_{T}=\sum_{t=1}^{T}\frac{t}{\sum_{t=1}^{T}t}\hat{\lambda_{t}}$. Here $\hat{\lambda}_{t,s}^{*}$ is calculated by the quadratic programming problem 2 with stochastic gradients. (Yang et al. 2024) uses $\Delta_{Q}^{t}=\sum_{s\in[S]}\lambda_{t,s}^{*}\left[f_{s}({\mathbf{x}}_{t})-f_{s}({\mathbf{x}}_{*})\right]$ as the metrics. For non-convex functions, $\|\mathbf{d}_{t}\|^{2}=\|\bm{\lambda}_{t}^{*T}\nabla\mathbf{F}({\mathbf{x}}_{t})\|^{2}$ (or equivalently $\mathbb{E}[\min_{\bm{\lambda}_{t}}\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]$) is usually used as the metrics. Our convergence metrics follow the conditions from  (Zhou et al. 2022). For non-convex functions, we use $\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]$ as the metrics, which can directly connect with metrics used in other works as $\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\min_{\bm{\lambda}_{t}}\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]\leq\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]$. The same applies for convex case as $\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\max_{{\mathbf{x}}_{t}^{*}}\min_{\bm{\lambda}_{t}^{*}}\left(G({\mathbf{x}}_{t},\bm{\lambda}_{t}^{*})-G({\mathbf{x}}_{t}^{*},\bm{\lambda}_{t}^{*})\right)]\leq\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[G({\mathbf{x}}_{t},\bm{\lambda}_{t})-G({\mathbf{x}}_{*},\bm{\lambda}_{t})]$.
Convergence Rates. As shown in Table 1, we compare the convergence rates with baselines across strongly-convex, general convex and non-convex functions. We show our PSMGD can achieve the state-of-the-art convergence rates in these cases, which match existing best results.
A.2 Auxiliary Lemma
Lemma A.1 (Lemma 2 and Lemma 7 in  (Zhou et al. 2022)).
When we calculate the $\hat{\bm{\lambda}}_{t}^{*}$ by QP using stochastic gradient $\nabla f_{s}({\mathbf{x}}_{t},\xi_{t})$, we have $\mathbb{E}_{\xi}\left<\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t}),-\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\right>\leq 2SBH\sqrt{m\sigma^{2}\mathbb{V}_{\xi}[\bm{\lambda}_{t}]}-\mathbb{E}_{\xi}[\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]$ and $\mathbb{V}_{\xi}[\bm{\lambda}_{t}]:=\mathbb{E}_{\xi}[\|\bm{\lambda}_{t}-\mathbb{E}[\bm{\lambda}_{t}]\|^{2}]\leq S^{2}B^{2}(1-\alpha_{t})^{2}.$
Proof.
$\mathbb{V}_{\xi}[\bm{\lambda}_{t}]$ $=\mathbb{E}_{\xi}[\|\bm{\lambda}_{t}-\mathbb{E}[\bm{\lambda}_{t}]\|^{2}]$ (9)
$\leq\mathbb{E}_{\xi}[\|\bm{\lambda}_{t}-\bm{\lambda}_{t-R}\|^{2}]$ (10)
$=\mathbb{E}_{\xi}[\|(1-\alpha_{t})(\hat{\bm{\lambda}}_{t}^{*}-\bm{\lambda}_{t-R})\|^{2}]$ (11)
$\leq(1-\alpha_{t})^{2}\mathbb{E}_{\xi}[\|(\hat{\bm{\lambda}}_{t}^{*}-\bm{\lambda}_{t-R})\|^{2}]$ (12)
$\leq(1-\alpha_{t})^{2}S^{2}B^{2}$ (13)
where the first inequality is due to $\mathbb{E}_{\xi}[\|\bm{\lambda}_{t}-\mathbb{E}[\bm{\lambda}_{t}]\|^{2}]\leq\mathbb{E}_{\xi}[\|\bm{\lambda}_{t}-a\|^{2}]$ for any $a$ and the update of $\bm{\lambda}$ is recomputed every $R$ iterations; and the last inequality follows from the fact $0\leq\hat{\bm{\lambda}}_{t,s}^{*},\bm{\lambda}_{t,s}\leq B$.
$\mathbb{E}_{\xi}\left<\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t}),-\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\right>$ (14)
$=\mathbb{E}_{\xi}\left<\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t}),-\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})+\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\right>$ (15)
$=\mathbb{E}_{\xi}\left<\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t}),-\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})+\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\right>$ (16)
$-\mathbb{E}_{\xi}[\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]$ (17)
$\mathbb{E}_{\xi}\left<\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t}),-\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})+\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\right>$ (18)
$=\mathbb{E}_{\xi}\left<\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t}),\sum_{s\in[S]}\bm{\lambda}_{t,s}(\nabla f_{s}({\mathbf{x}}_{t})-\nabla f_{s}({\mathbf{x}}_{t},\xi_{t}))\right>$ (19)
$=\underbrace{\mathbb{E}_{\xi}\left<\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t}),\sum_{s\in[S]}(\bm{\lambda}_{t,s}-\mathbb{E}_{\xi}[\bm{\lambda}_{t,s}])(\nabla f_{s}({\mathbf{x}}_{t})-\nabla f_{s}({\mathbf{x}}_{t},\xi_{t}))\right>}_{A_{1}}$ (20)
$\quad+\underbrace{\mathbb{E}_{\xi}\left<\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t}-\mathbb{E}_{\xi}[\bm{\lambda}_{t}]),\sum_{s\in[S]}\mathbb{E}_{\xi}[\bm{\lambda}_{t,s}](\nabla f_{s}({\mathbf{x}}_{t})-\nabla f_{s}({\mathbf{x}}_{t},\xi_{t}))\right>}_{A_{2}}$ (21)
$\quad+\underbrace{\mathbb{E}_{\xi}\left<\nabla\mathbf{G}({\mathbf{x}}_{t},\mathbb{E}_{\xi}[\bm{\lambda}_{t}]),\sum_{s\in[S]}\mathbb{E}_{\xi}[\bm{\lambda}_{t,s}](\nabla f_{s}({\mathbf{x}}_{t})-\nabla f_{s}({\mathbf{x}}_{t},\xi_{t}))\right>}_{A_{3}}$ (22)
$A_{1}=\mathbb{E}_{\xi}\left<\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t}),\sum_{s\in[S]}(\bm{\lambda}_{t,s}-\mathbb{E}_{\xi}[\bm{\lambda}_{t,s}])(\nabla f_{s}({\mathbf{x}}_{t})-\nabla f_{s}({\mathbf{x}}_{t},\xi_{t}))\right>$ (23)
$\leq\mathbb{E}_{\xi}\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|\mathbb{E}_{\xi}\left\|\sum_{s\in[S]}(\bm{\lambda}_{t,s}-\mathbb{E}_{\xi}[\bm{\lambda}_{t,s}])(\nabla f_{s}({\mathbf{x}}_{t})-\nabla f_{s}({\mathbf{x}}_{t},\xi_{t}))\right\|$ (24)
$\leq SBH\mathbb{E}_{\xi}\left[\sum_{s\in[S]}|\bm{\lambda}_{t,s}-\mathbb{E}_{\xi}[\bm{\lambda}_{t,s}]|\left\|(\nabla f_{s}({\mathbf{x}}_{t})-\nabla f_{s}({\mathbf{x}}_{t},\xi_{t}))\right\|\right]$ (25)
$\leq\frac{SBH}{2\epsilon}\sum_{s\in[S]}\mathbb{E}_{\xi}\left[|\bm{\lambda}_{t,s}-\mathbb{E}_{\xi}[\bm{\lambda}_{t,s}]|^{2}\right]$ (26)
$+\frac{\epsilon SBH}{2}\sum_{s\in[S]}\mathbb{E}_{\xi}\left[\left\|(\nabla f_{s}({\mathbf{x}}_{t})-\nabla f_{s}({\mathbf{x}}_{t},\xi_{t}))\right\|^{2}\right]$ (27)
$\leq\frac{SBH}{2\epsilon}\mathbb{V}_{\xi}[\bm{\lambda}_{t}]+\frac{\epsilon S^{2}BH}{2}\sigma^{2}$ (28)
$\leq SBH\sqrt{S\sigma^{2}\mathbb{V}_{\xi}[\bm{\lambda}_{t}]},$ (29)
where the last inequality follows from the fact $\epsilon=\frac{\sqrt{\mathbb{V}_{\xi}[\bm{\lambda}_{t}]}}{\sqrt{S\sigma^{2}}}$.
Similarly, using the same trick, we have the upper bound for $A_{2}$:
$A_{2}\leq SBH\sqrt{m\sigma^{2}\mathbb{V}_{\xi}[\bm{\lambda}_{t}]}.$ (30)
For $A_{3}$, it is easy to verify that $A_{3}=0$ as the randomness is only on the stochastic gradient which is unbiased estimations.
Combining $A_{1},A_{2}$ and $A_{3}$, we have
$\mathbb{E}_{\xi}\left<\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t}),-\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\right>$ (31)
$\leq 2SBH\sqrt{m\sigma^{2}\mathbb{V}_{\xi}[\bm{\lambda}_{t}]}-\mathbb{E}_{\xi}[\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]$ (32)
Lemma A.2 (Lemma 1 in  (Zhou et al. 2022)).
When we calculate the $\hat{\bm{\lambda}}_{t}^{*}$ by QP using stochastic gradient $\nabla f_{s}({\mathbf{x}}_{t},\xi_{t})$, we have $\|\mathbb{E}[\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})]-\mathbb{E}[\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})]\|^{2}\leq S\sigma^{2}\mathbb{V}_{\xi}[\bm{\lambda}_{t}]\leq\sigma^{2}S^{3}B^{2}(1-\alpha_{t})^{2}.$
A.3 Proof for Non-convex Functions
See 3.5
Proof.
Case I: When we calculate the $\hat{\bm{\lambda}}_{t}^{*}$ in one step ($t\%R==0$), we have the following due to $L-$smoothness:
$\mathbb{E}_{\xi}[\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t})]-\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})$ (33)
$\leq\mathbb{E}_{\xi}\left<\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t}),-\eta_{t}\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\right>$ (34)
$+\frac{1}{2}L\eta^{2}\mathbb{E}_{\xi}[\|\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]$ (35)
$\leq 2SBH\eta_{t}\sqrt{m\sigma^{2}\mathbb{V}_{\xi}[\bm{\lambda}_{t}]}-\eta_{t}\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}$ (36)
$+\frac{1}{2}L\eta_{t}^{2}\mathbb{E}_{\xi}[\|\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]$ (37)
$\mathbb{E}_{\xi}[\|\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]$ (38)
$=2\mathbb{E}_{\xi}[\|\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]+2\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}$ (39)
$\leq 2\mathbb{E}_{\xi}[\|\sum_{s\in[S]}\lambda_{t}^{s}(\nabla f_{s}({\mathbf{x}}_{t},\xi_{t})-\nabla f_{s}({\mathbf{x}}_{t}))\|^{2}]+2\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}$ (40)
$\leq 2\mathbb{E}_{\xi}\left[\left(\sum_{s\in[S]}\lambda_{t}^{s}\|(\nabla f_{s}({\mathbf{x}}_{t},\xi_{t})-\nabla f_{s}({\mathbf{x}}_{t})\|\right)^{2}\right]+2\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}$ (41)
$\leq 2S^{2}B^{2}\sigma^{2}+2\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}$ (42)
$\mathbb{E}_{\xi}[\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t})]-\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})$ (43)
$\leq\mathbb{E}_{\xi}\left<\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t}),-\eta_{t}\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\right>+\frac{1}{2}L\eta^{2}\mathbb{E}_{\xi}[\|\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]$ (44)
$=2SBH\eta_{t}\sqrt{S\sigma^{2}\mathbb{V}_{\xi}[\bm{\lambda}_{t}]}-\eta_{t}(1-L\eta_{t})\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}+L\eta_{t}^{2}S^{2}B^{2}\sigma^{2}$ (45)
$\leq 2S^{3/2}B^{2}H(1-\alpha_{t})\eta_{t}\sigma-\eta_{t}(1-L\eta_{t})\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}+L\eta_{t}^{2}S^{2}B^{2}\sigma^{2}$ (46)
If $\frac{1}{2L}\leq\eta_{t}<\frac{1}{L}$, then
$\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}$ (47)
$\leq\frac{2[\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbb{E}_{\xi}[\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t})]]}{\eta_{t}}$ (48)
$+\frac{4S^{3/2}B^{2}H(1-\alpha_{t})\eta_{t}\sigma+2L\eta_{t}^{2}S^{2}B^{2}\sigma^{2}}{\eta_{t}}$ (49)
Case II: When we reuse the coefficients from previous steps $\bm{\lambda}_{t}=\bm{\lambda}_{t-\tau}$ ($(t-\tau)\%R==0,t\%R\neq 0,\tau<R$), we have the following:
$\mathbb{E}_{\xi}[\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t})]-\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})$ (50)
$\leq\mathbb{E}_{\xi}\left<\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t}),-\eta_{t}\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\right>+\frac{1}{2}L\eta^{2}\mathbb{E}_{\xi}[\|\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]$ (51)
$\leq-\eta_{t}\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}+\frac{1}{2}L\eta_{t}^{2}\mathbb{E}_{\xi}[\|\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]$ (52)
where the last inequality follows from the fact that $\mathbb{E}_{\xi}[\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})]=\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})$ as a fixed $\bm{\lambda}_{t}$ is used and independent of current step sampling.
Similarly, if $\frac{1}{2L}\leq\eta_{t}<\frac{1}{L}$, then
$\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}$ $\leq\frac{2[\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbb{E}_{\xi}[\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t})]]+2L\eta_{t}^{2}S^{2}B^{2}\sigma^{2}}{\eta_{t}}$ (53)
For non-increasing learning rate $\eta_{t}$,
$\sum_{t\in[T]}\frac{\mathbb{E}[\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t})]}{\eta_{t}}$ (54)
$=\sum_{t\%R==0}\frac{\mathbb{E}[\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{t+R},\bm{\lambda}_{t})]}{\eta_{t}}$ (55)
$=\frac{\mathbb{E}[\mathbf{G}({\mathbf{x}}_{0},\bm{\lambda}_{0})-\mathbf{G}({\mathbf{x}}_{T},\bm{\lambda}_{T})]}{\eta_{T}}+\sum_{t\%R==0,t\neq 0}\frac{\mathbb{E}[\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t}-\bm{\lambda}_{t-R})]}{\eta_{t}}$ (56)
$=\frac{\mathbb{E}[\mathbf{G}({\mathbf{x}}_{0},\bm{\lambda}_{0})-\mathbf{G}({\mathbf{x}}_{T},\bm{\lambda}_{T})]}{\eta_{T}}$ (57)
$+\sum_{t\%R==0,t\neq 0}\frac{\mathbb{E}[\mathbf{G}({\mathbf{x}}_{t},(1-\alpha_{t})(\hat{\bm{\lambda}}_{t}-\bm{\lambda}_{t-R}))]}{\eta_{t}}$ (58)
$\leq\frac{2SBF}{\eta_{T}}+F\sum_{t\%R==0,t\neq 0}\mathbb{E}\left[\sum_{s\in[S]}(1-\alpha_{t})|\hat{\lambda}_{t}^{s}-\lambda_{t-R}^{s}|\right]$ (59)
where the last inequality is due to $|f_{s}({\mathbf{x}})|\leq F$.
Telescoping and rearranging,
$\frac{1}{T}\sum_{t\in[T]}\mathbb{E}\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}$ (60)
$\leq\frac{1}{T}\sum_{t\in[T]}\frac{2[\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbb{E}_{\xi}[\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t})]]+2L\eta_{t}^{2}S^{2}B^{2}\sigma^{2}}{\eta_{t}}$ (61)
$+\frac{4S^{3/2}B^{2}H\sigma}{T}\sum_{t\%R==0}(1-\alpha_{t})$ (62)
$\leq\frac{4SBF}{T\eta_{T}}+\frac{2F}{T}\sum_{t\%R==0,t\neq 0}\frac{\mathbb{E}\left[\sum_{s\in[S]}(1-\alpha_{t})|\hat{\lambda}_{t}^{s}-\lambda_{t-R}^{s}|\right]}{\eta_{t}}$ (63)
$\quad+2LS^{2}B^{2}\sigma^{2}\frac{1}{T}\sum_{t\in[T]}\eta_{t}+\frac{4S^{3/2}B^{2}H\sigma}{T}\sum_{t\%R==0}(1-\alpha_{t})$ (64)
By setting $1-\alpha_{t}=\min\{\frac{\eta_{t}}{\eta_{1}},\frac{\eta_{t}}{\eta_{1}\sqrt{t}\max_{s\in[S]}|\hat{\lambda}_{t}^{s}-\lambda_{t-R}^{s}|}\},$ and $\eta_{t}=\mathcal{O}(\frac{1}{\sqrt{t}})$, the convergence rate is $\frac{1}{T}\sum_{t\in[T]}\mathbb{E}\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}=\mathcal{O}(\frac{1}{\sqrt{T}}).$
A.4 Proof for General Convex Functions
See 3.6
Proof.
$\mathbb{E}[\|{\mathbf{x}}_{t+1}-{\mathbf{x}}_{*}\|^{2}]=\mathbb{E}[\|{\mathbf{x}}_{t}-\eta_{t}\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-{\mathbf{x}}_{*}\|^{2}]$ (65)
$=\|{\mathbf{x}}_{t}-{\mathbf{x}}_{*}\|^{2}+\eta_{t}^{2}\mathbb{E}[\|\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]$ (66)
$-2\eta_{t}\mathbb{E}\left<\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t}),({\mathbf{x}}_{t}-{\mathbf{x}}_{*})\right>$ (67)
$=\|{\mathbf{x}}_{t}-{\mathbf{x}}_{*}\|^{2}+\eta_{t}^{2}\mathbb{E}[\|\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]$ (68)
$-2\eta_{t}\left<\mathbb{E}[\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})]-\mathbb{E}[\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})],({\mathbf{x}}_{t}-{\mathbf{x}}_{*})\right>$ (69)
$\quad-2\eta_{t}\left<\mathbb{E}[\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})],({\mathbf{x}}_{t}-{\mathbf{x}}_{*})\right>$ (70)
Due to convexity, we have $\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t}),({\mathbf{x}}_{t}-{\mathbf{x}}_{*})\right>\geq\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{*},\bm{\lambda}_{t})$.
Case I: When we calculate the $\hat{\bm{\lambda}}_{t}^{*}$ in one step ($t\%R==0$), $\mathbb{E}[\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})]-\mathbb{E}[\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})]\neq 0$ due to dependence of $\bm{\lambda}_{t}$ on the samples.
$2\eta_{t}[\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{*},\bm{\lambda}_{t})]\leq-\mathbb{E}[\|{\mathbf{x}}_{t+1}-{\mathbf{x}}_{*}\|^{2}]+\|{\mathbf{x}}_{t}-{\mathbf{x}}_{*}\|^{2}$ (71)
$+\eta_{t}^{2}\mathbb{E}[\|\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]$ (72)
$-2\eta_{t}\left<\mathbb{E}[\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})]-\mathbb{E}[\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})],({\mathbf{x}}_{t}-{\mathbf{x}}_{*})\right>$ (73)
$\leq-\mathbb{E}[\|{\mathbf{x}}_{t+1}-{\mathbf{x}}_{*}\|^{2}]+\|{\mathbf{x}}_{t}-{\mathbf{x}}_{*}\|^{2}+\eta_{t}^{2}\mathbb{E}[\|\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]$ (74)
$+2\eta_{t}\|\mathbb{E}[\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})]-\mathbb{E}[\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})]\|\|({\mathbf{x}}_{t}-{\mathbf{x}}_{*})\|$ (75)
$\leq-\mathbb{E}[\|{\mathbf{x}}_{t+1}-{\mathbf{x}}_{*}\|^{2}]+\|{\mathbf{x}}_{t}-{\mathbf{x}}_{*}\|^{2}+2S^{2}B^{2}\eta_{t}^{2}\sigma^{2}$ (76)
$+2\eta_{t}^{2}\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}+2\eta_{t}D\sigma S^{3/2}B(1-\alpha_{t})$ (77)
$\leq-\mathbb{E}[\|{\mathbf{x}}_{t+1}-{\mathbf{x}}_{*}\|^{2}]+\|{\mathbf{x}}_{t}-{\mathbf{x}}_{*}\|^{2}+2S^{2}B^{2}\eta_{t}^{2}\sigma^{2}+2\eta_{t}^{2}S^{2}B^{2}H^{2}$ (78)
$+2\eta_{t}D\sigma S^{3/2}B(1-\alpha_{t})$ (79)
$\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{*},\bm{\lambda}_{t})\leq\frac{-\mathbb{E}[\|{\mathbf{x}}_{t+1}-{\mathbf{x}}_{*}\|^{2}]+\|{\mathbf{x}}_{t}-{\mathbf{x}}_{*}\|^{2}}{2\eta_{t}}$ (80)
$+\eta_{t}S^{2}B^{2}\sigma^{2}+\eta_{t}S^{2}B^{2}H^{2}+D\sigma S^{3/2}B(1-\alpha_{t})$ (81)
Case II: When we reuse the coefficients from previous steps $\bm{\lambda}_{t}=\bm{\lambda}_{t-\tau}$ ($(t-\tau)\%R==0,t\%R\neq 0,\tau<R$), $\mathbb{E}[\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})]-\mathbb{E}[\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})]=0$.
$2\eta_{t}[\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{*},\bm{\lambda}_{t})]$ (82)
$\leq-\mathbb{E}[\|{\mathbf{x}}_{t+1}-{\mathbf{x}}_{*}\|^{2}]+\|{\mathbf{x}}_{t}-{\mathbf{x}}_{*}\|^{2}+\eta_{t}^{2}\mathbb{E}[\|\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}]$ (83)
$\leq-\mathbb{E}[\|{\mathbf{x}}_{t+1}-{\mathbf{x}}_{*}\|^{2}]+\|{\mathbf{x}}_{t}-{\mathbf{x}}_{*}\|^{2}+2S^{2}B^{2}\eta_{t}^{2}\sigma^{2}+2\eta_{t}^{2}S^{2}B^{2}H^{2}$ (84)
That is,
$\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{*},\bm{\lambda}_{t})$ (85)
$\leq\frac{-\mathbb{E}[\|{\mathbf{x}}_{t+1}-{\mathbf{x}}_{*}\|^{2}]+\|{\mathbf{x}}_{t}-{\mathbf{x}}_{*}\|^{2}}{2\eta_{t}}+\eta_{t}S^{2}B^{2}\sigma^{2}+\eta_{t}S^{2}B^{2}H^{2}$ (86)
Assume $\eta_{t}=\frac{1}{\sqrt{t}}$ and $(1-\alpha_{t})=\eta_{t}$, combining two cases and telescoping:
$\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{*},\bm{\lambda}_{t})]$ (87)
$\leq\frac{\|{\mathbf{x}}_{1}-{\mathbf{x}}_{*}\|^{2}}{T\eta_{1}}+(2S^{2}B^{2}\sigma^{2}+2S^{2}B^{2}H^{2})\frac{1}{T}\sum_{t\in[T]}\eta_{t}$ (88)
$+2D\sigma S^{3/2}B\frac{1}{T}\sum_{t\%R==0}(1-\alpha_{t})$ (89)
$=\mathcal{O}(1/\sqrt{T})$ (90)
A.5 Proof for Strongly-Convex Functions
See 3.7
Proof.
Case I: When we calculate the $\hat{\bm{\lambda}}_{t}^{*}$ in one step ($t\%R==0$), $\mathbb{E}[\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})]-\mathbb{E}[\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})]\neq 0$ due to dependence of $\bm{\lambda}_{t}$ on the samples. Due to the strongly-convex property of $f_{s},s\in[S]$,
$2\mu\mathbb{E}_{\xi}(\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{t}^{*},\bm{\lambda}_{t}))\leq\mathbb{E}_{\xi}\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}$ (91)
$\leq\frac{2[\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbb{E}_{\xi}[\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t})]]+4S^{3/2}B^{2}H(1-\alpha_{t})\eta_{t}\sigma}{\eta_{t}}$ (92)
$+\frac{2L\eta_{t}^{2}S^{2}B^{2}\sigma^{2}}{\eta_{t}}$ (93)
The second inequality is due to  equation 49, where ${\mathbf{x}}_{t}^{*}=\operatorname*{arg\,min}_{{\mathbf{x}}}\mathbf{G}({\mathbf{x}},\bm{\lambda}_{t})$.
Rearranging it, we have
$\mathbb{E}_{\xi}[\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{t}^{*},\bm{\lambda}_{t})]$ (94)
$\leq(1-2\mu\eta_{t})\mathbb{E}_{\xi}(\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{t}^{*},\bm{\lambda}_{t}))$ (95)
$+4S^{3/2}B^{2}H(1-\alpha_{t})\eta_{t}\sigma+2L\eta_{t}^{2}S^{2}B^{2}\sigma^{2}.$ (96)
We know that the $\bm{\lambda}$ will be reused for $t+1\%\neq 0$. That is, $\bm{\lambda}_{t+1}=\bm{\lambda}_{t}$ and ${\mathbf{x}}_{t+1}^{*}={\mathbf{x}}_{t}^{*}$.
$\mathbb{E}_{\xi}[\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t+1})-\mathbf{G}({\mathbf{x}}_{t+1}^{*},\bm{\lambda}_{t+1})]$ (97)
$\leq(1-2\mu\eta_{t})\mathbb{E}_{\xi}(\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{t}^{*},\bm{\lambda}_{t}))$ (98)
$+4S^{3/2}B^{2}H(1-\alpha_{t})\eta_{t}\sigma+2L\eta_{t}^{2}S^{2}B^{2}\sigma^{2}.$ (99)
Case II: When we reuse the coefficients from previous steps $\bm{\lambda}_{t}=\bm{\lambda}_{t-\tau}$ ($(t-\tau)\%R==0,t\%R\neq 0,\tau<R$), $\mathbb{E}[\hat{\nabla}\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})]=\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})$.
$2\mu\mathbb{E}_{\xi}(\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{t}^{*},\bm{\lambda}_{t}))\leq\mathbb{E}_{\xi}\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}$ (100)
$\leq\frac{2[\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbb{E}_{\xi}[\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t})]]+2L\eta_{t}^{2}S^{2}B^{2}\sigma^{2}}{\eta_{t}}$ (101)
The last inequality follows from the fact that
$\|\nabla\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})\|^{2}$ (102)
$\leq\frac{2[\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbb{E}_{\xi}[\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t})]]+2L\eta_{t}^{2}S^{2}B^{2}\sigma^{2}}{\eta_{t}}$ (103)
Rearranging it, we have
$\mathbb{E}_{\xi}[\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{t}^{*},\bm{\lambda}_{t})]$ (104)
$\leq(1-2\mu\eta_{t})\mathbb{E}_{\xi}(\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{t}^{*},\bm{\lambda}_{t}))+2L\eta_{t}^{2}S^{2}B^{2}\sigma^{2}$ (105)
For $t+1\%R\neq 0$, we know that $\bm{\lambda}_{t+1}=\bm{\lambda}_{t}$ and ${\mathbf{x}}_{t+1}^{*}={\mathbf{x}}_{t}^{*}$ by reusing $\bm{\lambda}$, then
$\mathbb{E}_{\xi}[\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t+1})-\mathbf{G}({\mathbf{x}}_{t}^{*},\bm{\lambda}_{t+1})]$ (106)
$\leq(1-2\mu\eta_{t})\mathbb{E}_{\xi}(\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{t}^{*},\bm{\lambda}_{t}))+2L\eta_{t}^{2}S^{2}B^{2}\sigma^{2}$ (107)
For $t+1\%R=0$, we will recalculate $\bm{\lambda}_{t+1}$.
$\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t+1})-\mathbf{G}({\mathbf{x}}_{t+1}^{*},\bm{\lambda}_{t+1})$ (108)
$=\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{t+1}^{*},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t}-\bm{\lambda}_{t+1})$ (109)
$+\mathbf{G}({\mathbf{x}}_{t+1}^{*},\bm{\lambda}_{t}-\bm{\lambda}_{t+1})$ (110)
$\leq\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{t+1}^{*},\bm{\lambda}_{t})+2F(1-\alpha_{t})\sum_{s\in[S]}|\hat{\lambda}_{t+1}^{s}-\lambda_{t}^{s}|$ (111)
$\leq(1-2\mu\eta_{t})\mathbb{E}_{\xi}(\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{t}^{*},\bm{\lambda}_{t}))$ (112)
$+2L\eta_{t}^{2}S^{2}B^{2}\sigma^{2}+2F(1-\alpha_{t})\sum_{s\in[S]}|\hat{\lambda}_{t+1}^{s}-\lambda_{t}^{s}|$ (113)
$\leq(1-2\mu\eta_{t})\mathbb{E}_{\xi}(\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{t}^{*},\bm{\lambda}_{t}))+2L\eta_{t}^{2}S^{2}B^{2}\sigma^{2}$ (114)
$+4F(1-\alpha_{t})SB$ (115)
$\mathbf{G}({\mathbf{x}}_{t+1},\bm{\lambda}_{t+1})-\mathbf{G}({\mathbf{x}}_{t+1}^{*},\bm{\lambda}_{t+1})$ (116)
$\leq(1-2\mu\eta_{t})\mathbb{E}_{\xi}(\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{t}^{*},\bm{\lambda}_{t}))+2L\eta_{t}^{2}S^{2}B^{2}\sigma^{2}$ (117)
$\quad+4F(1-\alpha_{t})SB\mathbf{1}\{t+1\%R=0\}$ (118)
$\quad+4S^{3/2}B^{2}H(1-\alpha_{t})\eta_{t}\sigma\mathbf{1}\{t\%R=0\}$ (119)
$\quad\leq(1-2\mu\eta_{t})\mathbb{E}_{\xi}(\mathbf{G}({\mathbf{x}}_{t},\bm{\lambda}_{t})-\mathbf{G}({\mathbf{x}}_{t}^{*},\bm{\lambda}_{t}))$ (120)
$\quad+\eta_{t}^{2}(2LS^{2}B^{2}\sigma^{2}+4FSB\mathbf{1}\{t+1\%R=0\}$ (121)
$\quad+4S^{3/2}B^{2}H\sigma\mathbf{1}\{t\%R=0\})$ (122)
$\Phi=(2LS^{2}B^{2}\sigma^{2}+4FSB\mathbf{1}\{t+1\%R=0\}$ (123)
$+4S^{3/2}B^{2}H\sigma\mathbf{1}\{t\%R=0\}$ (124)
where we set $(1-\alpha_{t})=\eta_{t}$.
Using $\eta_{t}=\frac{c}{T}$ with $c>\frac{1}{2\mu}$,
$\mathbf{G}({\mathbf{x}}_{T},\bm{\lambda}_{T})-\mathbf{G}({\mathbf{x}}_{T}^{*},\bm{\lambda}_{T})$ (125)
$\leq\frac{\max\{2c^{2}\Phi(2\mu c-1)^{-1},\mathbf{G}({\mathbf{x}}_{0},\bm{\lambda}_{0})-\mathbf{G}({\mathbf{x}}_{0}^{*},\bm{\lambda}_{0})\}}{T}$ (126)
Appendix B Supplement to Experiments
B.1 Supplement convergence experiments for the QM9 and NYU-v2
Here are convergence experiments results for the QM9 and NYU-v2. Figure 2 illustrates the test loss trajectories over 300 training epochs for the QM-9 dataset, where our PSMGD is tasked with predicting various molecular properties. The figure compares the performance of PSMGD with other MOO algorithms. The test loss in Figure 2, computed as the average MAE across all molecular property prediction tasks, provides a clear metric of the convergence performance. The results demonstrate that PSMGD achieves a smoother and faster reduction in test loss, particularly during the early stages of training, showcasing its efficiency and robustness in handling the MOO challenges presented by the QM-9 dataset.
Figure 2: Test loss in training (300 epochs) on QM-9.
Figure 3 presents the test loss curves across 200 training epochs for three dense prediction tasks, including image segmentation, depth prediction, and surface normal estimation on the NYU-v2 dataset. It provides a comparative analysis of PSMGD against other gradient-based multi-objective optimization methods, highlighting its superior performance in reducing test loss across all tasks. The consistent decrease in loss for each task, particularly in the early training epochs, shows PSMGD’s ability to effectively manage and optimize multiple objectives simultaneously, making it a excellent algorithm for complex MOO scenarios such as those encountered in dense prediction datasets like the NYU-v2. Through the realistic convergence experiments here, we can clearly find PSMGD can achieve an outstanding convergence performance with advanced performance metrics in all tasks.
Figure 3: Test loss of segmentation, depth, and surface normal in training (200 epochs) on NYU-v2.
B.2 Pre-Study: Multi-MNIST
We present experiments on the Multi-MNIST dataset (Sener and Koltun 2018), a classical 2-task supervised learning benchmark. It is constructed by uniformly sampling MNIST (LeCun et al. 1998) images and placing one in the top-left corner and the other in the bottom-right corner. To ensure consistency with the experimental preparation of (Kurin et al. 2022), we utilize a modified encoder-decoder version of the LeNet architecture (LeCun et al. 1998). The cross-entropy loss is utilized for both tasks. All methods are trained in 100 epochs using Adam (Kingma and Ba 2014) in the stochastic gradient setting, with an initial learning rate of $21e-$ 2. They are tuned in $\{1e-3,1e-2,1e-1\}$ and yield the best validation results for all optimizers under consideration, exponentially decayed by 0.95 after each epoch with a batch size of 256. Note that, in the Multi-MNIST, every method is repeated 10 times using varying random seeds with the mean results reported.
Figure 4 displays the average test accuracy and the training time per epoch and iteration. The test model was chosen in each run based on the highest average validation accuracy across the configured epochs. In Figure 4(b), the overlapping confidence intervals of average test accuracy indicate that none of the considered algorithms outperforms the others. Their test accuracies in 10 runs of the training process are between 0.935 and 0.95.
In Figure 4(b) and (c), we conclude the results of the time consumption per epoch and per iteration. They also indicate that PSMGD has among the lowest training times, both per epoch and per iteration. Heading further into a training epoch, we discover that linear sacralization requires the least amount of time. PSMGD takes $1.44s$, which is slightly longer than linear sacralization but shorter than the other methods. Additionally, we can see that PSMGD only takes $3.58ms$, the shortest time for a training iteration, and also second only to linear sacralization.
However, in Figure 5(b), linear scalarization requires the most epochs to reach convergence (validation accuracy $\geq$ 0.9), resulting in a longer time despite having only one backpropagation per iteration. PSMGD boasts minimal backpropagation complexity and convergence time due to its streamlined but efficient structure. In particular, in Figure 5(c), PSMGD takes $7.19s$, the shortest average convergence time consumption, to achieve convergence in a single run.
In Figure 5(a), we also observe that PSMGD has a wider accuracy distribution in both Task 1 and Task 2. We use linear scalarization with grid search to simulate the approximate curve of the Pareto front for the Multi-MNIST dataset. Here, PSMGD’s task accuracies exhibit a broader distribution, with a pattern that encompasses single-task accuracies adjacent to task 1 and task 2, as well as their overlapping center regions.
Moreover, we also explore the weight adaptation here to validate the motivation of our algorithm PSMGD. During the MGDA training process, weights are updated regularly. Figure 6(a) shows that the weights for task 1 and task 2 have a consistent average value and low variability. This indicates that the weight vectors are updated minimally throughout the training process. Delving into an epoch, in Figure 6(b), it is evident that most weight updates fall within the range of [0.4, 0.6] per iteration. This suggests that MGDA does not make drastic changes to the weight vectors. This is precisely why we can update the weight vectors every $R$ iteration.
Figure 4: PSMGD achieves a favorable balance between test accuracy and training time of epoch and iteration. (a) (Left) Average test accuracy: mean and 95% CI (10 runs). (b) (Middle) Box plots for the average training time of an epoch (10 runs). (c) (Right) Box plots for the average training time of an iteration (10 epochs).
Figure 5: PSMGD achieves a favorable balance between each task’s accuracy and convergence complexity time of epoch in MultiMNIST. (a) (Left) The results for MultiMNIST with Task1&2 accuracy. (b) (Middle) Average backpropagation complexity for convergence: mean and 95% CI (10 runs). (c) (Right) Box plots for convergence time (10 runs).
Figure 6: The dynamic weight vectors for MGDA and linear scalarization during the training process for the MultiMNIST. (a) (Upper) The results for dynamic weight vectors for MGDA in a run (100 epochs). (b) (Lower) The results for dynamic weight vectors for MGDA in an epoch (196 iterations).
B.3 Additional Supervised Learning Experiments: CityScapes and CelebA
Experiment Environments. 1. Hyper-parameters. For our PSMGD, we choose the best hyperparameter $R\in\{4,8,16\}$ and momentum $\bm{\lambda}=0.9$ based on the validation loss in experiments. Specifically, we choose $R=4$ for the Multi-MNIST dataset in Section B.2 and $R=8$ for the rest of the datasets in Section 4. The hyper-parameters in the baselines follow from the settings in their corresponding papers, which align with existing works. 2. Software and Hardware. Following experiment settings from (Liu et al. 2024; Ban and Ji 2024; Xiao, Ban, and Ji 2024), all experiments except the Multi-MNIST are conducted using PyTorch on a single NVIDIA A100 GPU. The Multi-MNIST experiment is conducted on a single NVIDIA GeForce RTX 2070 Super GPU under Ubuntu 18.04 LTS.
CityScapes dataset (Cordts et al. 2016) is a comprehensive collection of high-quality street-scene images, specifically designed for the evaluation of computer vision algorithms in urban environments. It comprises 5K finely annotated images captured across different cities. The dataset supports 2 tasks: semantic segmentation, which involves classifying each pixel in an image into one of 30 classes grouped into 7 super-categories, and depth estimation, which involves predicting the distance of objects from the camera. These 2 tasks provide a robust benchmark for evaluating the effectiveness of MOO methods in handling complex urban scenes, characterized by a diverse range of objects, varying lighting conditions, and dynamic elements like pedestrians and vehicles. We follow experiment settings from (Liu et al. 2024; Ban and Ji 2024; Xiao, Ban, and Ji 2024) and adopt the backbone of MTAN (Liu, Johns, and Davison 2019), adding objective-specific attention modules to SegNet (Badrinarayanan, Kendall, and Cipolla 2017). Our method is trained for 200 epochs with batch size 8 on the CityScapes. The initial learning rate is set to $41e-$ 4 for the first 100 epochs and then halved for the remaining epochs.
CelebA dataset (Liu et al. 2015) is a large-scale face attributes dataset, consisting of over 200K images of celebrities, annotated with 40 different attributes. These attributes encompass a wide range of facial features and expressions, such as smiling, wavy hair, mustache, and many more. Each image is labeled with binary indicators for the presence or absence of these attributes, making CelebA a valuable resource for MOO research. In this context, it can be viewed as an image-level 40-task MTL classification problem, where each task involves predicting the presence of a specific attribute. This setup provides a comprehensive benchmark for evaluating the performance of MTL methods in managing a substantial number of tasks simultaneously, thereby highlighting their ability to capture and leverage shared information across different attributes. We replicate the experiment setup from (Liu et al. 2024; Ban and Ji 2024). The model consists of a 9-layer convolutional neural network (CNN) as the backbone, with a dedicated linear layer for each objective. Training is performed over 15 epochs, utilizing the Adam optimizer (Kingma and Ba 2014) with a learning rate of $43e-$ 4, and a batch size of 256.
In Table 6, we observe that PSMGD demonstrates commendable performance in the final results of CityScapes and CelebA. In particular, in CelebA, while $\Delta m\%$ of PSMGD ranks 6th best, its MR ranks the highest among all listed MOO algorithms. This indicates that PSMGD exhibits a balanced capability in dealing with a significant number of objectives concurrently.
Method CityScapes CelebA
Segmentation Depth MR $\downarrow$ $\Delta m\%$downarrow MR $\downarrow$ $\Delta m\%$downarrow
mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$
STL 74.01 93.16 0.0125 27.77
LS 70.95 91.73 0.0161 33.83 10.00 22.60 7.55 4.15
SI 70.95 91.73 0.0161 33.83 12.50 14.11 9.15 7.20
RLW 74.57 93.41 0.0158 47.79 12.75 24.38 6.38 1.46
DWA 75.24 93.52 0.0160 44.37 9.75 21.45 8.25 3.20
UW 72.02 92.85 0.0140 30.13 8.75 5.89 6.80 3.23
MGDA 68.84 91.54 0.0309 33.50 13.00 44.14 12.80 14.85
PCGrad 75.13 93.48 0.0154 42.07 10.25 18.29 7.95 3.17
GradDrop 75.27 93.53 0.0157 47.54 9.00 23.73 9.30 3.29
CAGrad 75.16 93.48 0.0141 37.60 8.75 11.64 7.35 2.48
IMTL-G 75.33 93.49 0.0135 38.41 6.50 11.10 5.67 0.84
MoCo 75.42 93.55 0.0149 34.19 5.25 9.90 $\times$ $\times$
NashMTL 75.41 93.66 0.0129 35.02 3.50 6.82 5.97 2.84
FAMO 74.54 93.29 0.0145 32.59 9.00 8.13 5.75 1.21
FairGrad 75.72 93.68 0.0134 32.25 1.50 5.18 6.47 0.37
SDMGrad 74.53 93.52 0.0137 34.01 7.50 7.79 $\times$ $\times$
PSMGD 74.90 93.37 0.0135 35.78 8.50 8.76 5.60 2.60
Table 6: Results for the CityScapes (2 tasks) and CelebA (40 tasks) datasets. We run each experiment three time using random seeds with the mean reported. The best result is highlighted in bold.
B.4 Ablation on $R$
To better analyze the effect of the hyperparameter $R$, we present the ablation study on the QM-9 dataset. We choose $R\in\{4,8,16,32\}$ and momentum $\bm{\lambda}=0.9$ in averaged over 3 random seeds.
In Figure 7, it can be observed that the test loss increases with a higher hyperparameter $R$. Nevertheless, there is a minimal difference in the final convergences between PSMGD-4 and PSMGD-8, indicating proximity to the best convergence with $R=8$. For comparison, LS and MGDA are also included in Table 7. We note that the training time per epoch will rise as $R$ decreases, indicating that solving the optimization problem  2 will result in increased time consumption. Among the final results in $\Delta m\%$, PSMGD-4 demonstrates superior performance with the longest training time in PSMGD series algorithms. While PSMGD-16 and PSMGD-32 have a shorter training duration, their $\Delta m\%$ values do not match those of PSMGD-8.
Figure 7: Test loss in training (300 epochs) on QM-9.
Method Training time $\downarrow$ $\Delta m\%$downarrow
(mean per epoch)
LS 1.70 177.6
MGDA 12.94 120.5
PSMGD-4 2.61 88.2
PSMGD-8 2.23 92.4
PSMGD-16 2.13 109.7
PSMGD-32 2.07 116.2
Table 7: Training time per epoch [Min.] (averaged over 3 random seeds) on QM-9.

### v2

@@
 Abstract
-Multi-objective optimization (MOO) lies at the core of many machine learning (ML) applications that involve multiple, potentially conflicting objectives (e.g., multi-task learning, multi-objective reinforcement learning, among many others).
+Multi-objective optimization (MOO) lies at the core of many machine learning (ML) applications that involve multiple, potentially conflicting objectives.
 Despite the long history of MOO, recent years have witnessed a surge in interest within the ML community in the development of gradient manipulation algorithms for MOO, thanks to the availability of gradient information in many ML problems.
@@
 Codehttps://github.com/MarcuXu/PSMG
-Extended versionhttps://aaai.org/example
+Extended versionhttps://arxiv.org/abs/2412.10961
 1 Introduction
-Just as the human world is full of diverse and potentially conflicting goals, many machine learning paradigms are also multi-objective, such as multi-task learning (Caruana 1997; Sener and Koltun 2018), multi-objective reinforcement learning (Hayes et al.
-2022), learning-to-rank (Mahapatra et al.
-2023; Mahapatra, Dong, and Momma 2023), among others.
+Just as the human world is full of diverse and potentially conflicting goals, many machine learning paradigms are also multi-objective, such as multi-objective reinforcement learning (Hayes et al.
+2022), multi-task learning (Caruana 1997; Sener and Koltun 2018), and learning-to-rank (Mahapatra et al.
+2023).
 At the core of these machine learning paradigms lies multi-objective optimization (MOO), which aims to optimize multiple potentially conflicting objectives simultaneously.
@@
 where ${\mathbf{x}}\in\mathcal{D}\subseteq\mathbb{R}^{d}$ is the model parameter, and $f_{s}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, $s\in[S]$ is the objective function.
-The most common approach for handling MOO is to optimize a weighted average of the multiple objectives, also known as static/unitary weighting or linear scalarization (Kurin et al.
+The most common approach for handling MOO is to optimize a weighted average of the multiple objectives, also known as linear scalarization (Kurin et al.
 2022; Xin et al.
@@
 The linear scalarization approach has a fast per-iteration runtime but suffers from slow convergence and poor performance due to potential conflicts among objectives.
-To address this problem, there has been a surge of interest in recent years in developing gradient manipulation algorithms for MOO (Kendall, Gal, and Cipolla 2018; Chen et al.
+To address this problem, there has been a surge of interest in recent years in developing gradient manipulation algorithms (Kendall, Gal, and Cipolla 2018; Chen et al.
 2018; Yu et al.
 2020; Chen et al.
-2020; Javaloy and Valera 2021; Liu et al.
-2021a; Liu and Vicente 2021; Chen et al.
+2020; Javaloy and Valera 2021; Liu and Vicente 2021; Chen et al.
 2024; Xiao, Ban, and Ji 2024).
-The key idea is to calculate dynamic weights to avoid conflicts and find a direction $\mathbf{d}({\mathbf{x}})$ that decreases all objective functions simultaneously.
+The key idea is to calculate dynamic weights to avoid conflicts and find a direction $\mathbf{d}({\mathbf{x}})$ that decreases all objectives simultaneously.
 This common descent vector $\mathbf{d}({\mathbf{x}})$ is often determined by solving an additional optimization problem that involves all objective gradients.
@@
 To this end, a natural question arises: Can we design efficient gradient manipulation algorithms for fast MOO with theoretical guarantees?
-In this paper, we present Periodic Stochastic Multi-Gradient Descent (PSMGD), a simple yet effective algorithm to accelerate the MOO process.
+In this paper, we present Periodic Stochastic Multi-Gradient Descent (PSMGD), a simple yet effective algorithm to accelerate the MOO.
 Our PSMGD is motivated by the key observation that dynamic weights across objectives exhibit small changes under minor updates over short intervals.
-Thus, our PSMGD algorithm is designed to periodically compute these dynamic weights and utilizes them repeatedly, thereby effectively reducing the computational overload and achieving a faster MOO training process.
+Thus, it is designed to periodically compute these dynamic weights and utilizes them repeatedly, thereby effectively reducing the computational overload and achieving a faster MOO training process.
 Our contributions are summarized as follows:
@@
   • 2. Notations: T is the number of iterations, S is the number of objectives, R is the weight calculation hyper-parameter.
-  • Built upon the key observation, we introduce PSMGD, a new and efficient gradient manipulation method algorithm.
-By calculating dynamic weights across objectives infrequently, PSMGD significantly alleviates computational burden and reduces training time.
-  • We conduct a theoretical analysis of PSMGD for strongly convex, general convex, and non-convex functions, demonstrating that PSMGD achieves state-of-the-art convergence rates comparable to existing MOO methods (Table 1).
+  • Built upon the key observation, we introduce PSMGD, a new and efficient gradient manipulation method.
+By calculating dynamic weights infrequently, it significantly alleviates computational burden and reduces training time.
+  • We conduct a theoretical analysis of PSMGD for strongly convex, convex, and non-convex functions, demonstrating that PSMGD achieves state-of-the-art convergence rates comparable to existing MOO methods (Table 1).
 We also introduce a new computational complexity measure, backpropagation (BP) complexity, to quantify computational workload, and further show that PSMGD can achieve an objective-independent backpropagation complexity.
-Compared with existing MOO methods, this shows a superior BP complexity with the linear speedup with respect to the number of objectives.
   • Through comprehensive evaluation, PSMGD shows comparable or even superior performance compared to existing MOO methods, with significantly less training time.
@@
 The MOO problem, as stated in Eq.
-1, has a long and rich history that dates back to the 1950s.
+1, has a long history that dates back to the 1950s.
 MOO algorithms can be broadly categorized into two main groups.
@@
 2020; Laumanns and Ocenasek 2002).
-These methods are more suitable for small-scale problems but are less practical for high-dimensional MOO models, such as deep neural networks.
+These methods are more suitable for small-scale problems but are less practical for high-dimensional models, such as deep neural networks.
 The second category is the gradient-based approach by utilizing (stochastic) gradients (Fliege and Svaiter 2000; Désidéri 2012; Fliege, Vaz, and Vicente 2019; Liu and Vicente 2021), making them more practical for high-dimensional MOO problems.
 In this work, we primarily focus on gradient-based approaches to solve MOO in high-dimensional deep-learning models.
-In what follows, we provide a brief overview of two typical approaches for gradient-based methods for MOO: linear scalarization (LS) and gradient manipulation methods.
-1) LS: A straightforward approach for MOO 1 is to transform it into a single objective problem by using pre-defined weights $\{\lambda_{s},s\in[s]\}$: $\min_{{\mathbf{x}}\in\mathcal{D}}\sum_{s\in[S]}\lambda_{s}f_{s}({\mathbf{x}}).$ Thanks to the LS, one can leverage many existing and familiar single-objective methods, such as stochastic gradient descent, for training at each iteration.
+We first provide a brief overview of two typical approaches: linear scalarization (LS) and gradient manipulation methods.
+1) LS: A straightforward approach is to transform it into a single objective problem by using pre-defined weights $\bm{\lambda}$: $\min_{{\mathbf{x}}\in\mathcal{D}}\bm{\lambda}^{\top}\mathbf{F}({\mathbf{x}}).$ Thanks to the LS, one can leverage many existing single-objective methods (e.g., gradient descent), for training at each iteration.
 However, in LS, it is not uncommon to see conflicts among multiple objectives during the optimization process, i.e., $\nabla f_{s}({\mathbf{x}}),\nabla f_{s^{\prime}}({\mathbf{x}})\right><0$.
 This implies that the update using static weights may decrease some objectives, while inevitably increasing others at the same time and leading to slow convergence and poor performance.
-2) Gradient Manipulation Methods: A popular alternative is to dynamically weight gradients across objectives to avoid such conflicts and obtain a direction to decrease all objectives simultaneously.
-For example, multiple gradient descent algorithm (MGDA) (Fliege and Svaiter 2000) is seeking to find the common descent direction by solving the following optimization problem given each objective’s gradient: $(\mathbf{d},\beta)\in\operatorname*{arg\,min}_{\mathbf{d}\in\mathbb{R}^{S},\beta\in\mathbb{R}}\beta+\frac{1}{2}\|\mathbf{d}\|^{2},s.t.,\nabla f_{s}({\mathbf{x}})^{T}\mathbf{d}-\beta\leq 0,\forall s\in[S].$ By doing so, if ${\mathbf{x}}$ represents a first-order Pareto stationary point, then $(\mathbf{d},\beta)=(\mathbf{0},0)$.
+2) Gradient Manipulation Methods: A popular alternative is to dynamically weight gradients across objectives to avoid such conflicts and obtain a direction $\mathbf{d}$ to decrease all objectives simultaneously.
+For example, multiple gradient descent algorithm (MGDA) (Fliege and Svaiter 2000) seeks to find the $\mathbf{d}$ by solving the following optimization problem given gradients: $(\mathbf{d},\beta)\in\operatorname*{arg\,min}_{\mathbf{d}\in\mathbb{R}^{S},\beta\in\mathbb{R}}\beta+\frac{1}{2}\|\mathbf{d}\|^{2},s.t.,\nabla f_{s}({\mathbf{x}})^{T}\mathbf{d}-\beta\leq 0,\forall s\in[S].$ By doing so, if ${\mathbf{x}}$ represents a first-order Pareto stationary point, then $(\mathbf{d},\beta)=(\mathbf{0},0)$.
 Otherwise, we can find the $\mathbf{d}$ such that $\nabla f_{s}({\mathbf{x}})^{T}\mathbf{d}\leq\beta<0,s\in[S]$ as the solution.
@@
 However, due to the additional optimization process required to generate such $\mathbf{d}$, gradient manipulation methods often incur an expensive per-iteration cost, thereby prolonging training time.
-In this work, we propose a new algorithm, PSMGD, which combines the advantages of both approaches and offers a fast per-iteration runtime, a fast overall convergence speed, and improved model performance.
+In this work, we propose a new algorithm, PSMGD, which achieves fast per-iteration and overall convergence, along with enhanced performance.
 3 Periodic stochastic multi-gradient descent
@@
 We verified stability by examining two scales, epoch and iteration.
-The weights show consistent behavior with a stable average value and low variability across both scales, epoch and iteration, as depicted in Figure 6(a) and 6(b).
+The weights show consistent behavior with a stable average value and low variability across both scales, epoch and iteration, as depicted in Figure 6(a) and 6(b) in Appendix.
 This observation suggests a new approach to develop more efficient MOO algorithms.
@@
 3.4 Convergence analysis
-Before providing the convergence analysis, we have the following assumptions:
 Assumption 3.1 (L-Lipschitz continuous).
-There exists a constant $L>0$ such that $\|\nabla f_{s}({\mathbf{x}})-\nabla f_{s}({\mathbf{y}})\|\leq L\|{\mathbf{x}}-{\mathbf{y}}\|,\forall{\mathbf{x}},{\mathbf{y}}\in\mathbb{R}^{d},s\in[S]$.
-Assumption 3.2 (Unbiased stochastic gradient with bounded variance).
-We assume the stochastic gradient estimation of each objective function is unbiased with bounded variance.
+$\|\nabla f_{s}({\mathbf{x}})-\nabla f_{s}({\mathbf{y}})\|\leq L\|{\mathbf{x}}-{\mathbf{y}}\|,\forall{\mathbf{x}},{\mathbf{y}}\in\mathbb{R}^{d},s\in[S]$.
+Assumption 3.2 (Bounded variance).
+We assume the stochastic gradient estimation is unbiased with bounded variance.
 $\mathbb{E}\left[\nabla f_{s}({\mathbf{x}},\xi)\right]=\nabla f_{s}({\mathbf{x}}),$ (3)
@@
 We believe future work could explore preference-based solutions or the entire Pareto set using our efficient PSMGD algorithm.
+Acknowledgments
+JL acknowledges the funding from NSF grants CAREER CNS-2110259, CNS-2112471, IIS-2324052, DARPA YFA D24AP00265, ONR grant N00014-24-1-2729, and AFRL grant PGSC-SC-111374-19s.
+HY acknowledges the funding support from AI Seed Funding and GWBC Award at RIT.
 Appendix